What is the maximum height of the equation
y= -.015x^2 - 275x

Answer by Theo(13342)

since this quadratic is in standard form of y = ax^2 + bx + c, then a = -.015 and b = -275 and c = 0

the maximum height will be when x = -b/2a.

s = -b/2a becomes x = -275 / (2 * -.015) which becomes x = -275 / -.03 which becomes x = 9166.66666... = 9166 and 2/3.

the roots of this equation can be found using the quadratic formula.

the roots are x = -18333.33333.... and x = 0.

here's a graph of the equation with the maximum value and the value of the roots shown.

$$$